OpenStudy (anonymous):

writing a system of equations that has each ordered pair as its solution? (ex: 5,4)

OpenStudy (anonymous):

x=5 y=4 so make up two random expression such as x+y or 3x-4y then solve them both and put them in the form Ax+By=c

OpenStudy (anonymous):

needs to be 2 separate equations.

OpenStudy (anonymous):

x+y=? x-y=?

OpenStudy (anonymous):

yess..but how do i go from (5,4) to 2 different equations.

OpenStudy (anonymous):

answer the dang question and you'll find out

OpenStudy (anonymous):

all i am given is "write a system of linear equations that has a order pair of (5,4)" .

OpenStudy (anonymous):

given (5,4) meaning x= 5 y=4 x+y=? x-y=?

OpenStudy (anonymous):

uhh that's definitely not it. i know that for sure. but thanks..

OpenStudy (anonymous):

x+y=5+4=????? x-y=5-4=??????

OpenStudy (anonymous):

you're not being clear..haha

OpenStudy (anonymous):

UGGGGGGG ill telling showing you how would would create a system of equations......... first you have a coordinate point (5,4) coordinate points come in the form of (x,y) thus x=5 y=4 in order to get a system of equation with only these points as a solution you need to create 2 equations that when you substitute these values into the equation will make it true system of equations usually come in the form of Ax+By=C where A, B, and C are constants you can make up random values for A and B, C can solved when you substitute x and y into the equation example, if A=1 and B=1 then i will get x+y=C now i solve for C 5+4=C C=9 thus meaning one of the equations for the system of equations can be x+y=9 now you need to create a 2nd equation that is true when you substitute in x and y in this case, i chose A=1 and B=-1 x-y=C finish it......

OpenStudy (anonymous):

chill yourself. mahhh goodness. not my fault what you're saying doesn't make sense when, i know this much, theres gotta be numbers infront of x and/or y, not x+y equalying 9. get my drift?

OpenStudy (anonymous):

there are numbers infront of x and y 1 is a number if you dont like 1, then use 2 all i gave was an example